# thanks: These authors contributed equallythanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@
Atomically thin van der Waals (vdW) magnets have emerged as a fascinating platform for the exploration of novel physical phenomena arising from their reduced dimensionality and exceptional material properties. Their single-crystalline nature and ultimate miniaturization position them as leading candidates for next-generation spintronic applications. Antiferromagnetic (AF) vdW magnets are of particular interest, as they combine the advantages of vdW magnets with the functionality of AF spintronics, offering unique opportunities for ultrafast and robust spintronic devices. However, the lack of approaches to locally and deterministically manipulate their order parameter – the Néel-vector – remains a key limitation. Here, we introduce a fundamentally new paradigm in nanomagnetism, which we term lateral exchange bias (LEB), to achieve Néel vector control in bilayers of the vdW AF CrSBr. We exploit the single-crystalline registry formed by terraced CrSBr samples, where the bilayer Néel vector is controlled by LEB from neighboring, odd-layered flakes, whose nonzero magnetization we manipulate using magnetic fields. Using this control, we achieve nonvolatile manipulation of magnetic domains and domain walls (DWs) in AF CrSBr bilayers, establishing a powerful toolkit for controlling atomically thin AFs at the nanoscale. Our results challenge conventional views on exchange bias and provide a previously unexplored mechanism for achieving atomic-scale control of AFic order. Our findings pave the way for the development of advanced spintronic architectures and quantum technologies based on vdW magnets.
@@
The order parameter of few-layer vdW magnets can be controlled by a range of external stimuli, including magnetic 1, 2 and electric fields 3, 4, doping 4, 5, pressure 6, and strain 7. However, current approaches do not provide the sort of local control over order parameters that is required to deterministically create spin textures such as magnetic domain walls (DWs), and, in most cases, do not allow bistable switching between nonvolatile states which is a prerequisite for vdW spintronics. While important first steps have been taken in this direction 8, realizing such local control in fully compensated vdW AFs remains challenging, and so far no controlled writing of spin textures has been demonstrated.
@@
In classical magnetism 9, 10 and thin-film spintronics devices 11, 12, the exchange bias (EB) phenomenon has been known for decades as an efficient approach to controlling magnetic order parameter, and was instrumental for the development of memory and sensor technologies 13. The EB effect is based on interfacial exchange interactions between two adjacent magnetic layers, where the order parameter of a pinning layer, typically an AF, controls the order parameter in the adjacent target layer, typically a ferromagnet (FM).
Despite intense research on EB in vdW magnets 14, 15, 16, 17, its use for vdW spintronics and order-parameter control in vdW magnets remains challenging due to the weakness of the interlayer EB and difficulties in reproducibly controlling the vdW gap in vertical vdW heterostructures 18.
Here, we introduce a novel concept of EB in vdW AFs that is intrinsically free of these limitations and demonstrate its use for local control of the order parameter and spin-texture initialization in bilayers of the vdW AF CrSBr 19.
@@
Our concept, which we term lateral exchange bias (LEB), exploits exchange interactions across the near-perfect one-dimensional interface between two adjacent regions with different numbers of layers of the same single-crystal vdW magnet for order-parameter control. Lateral exchange has previously been observed with FM thin films 20, 21 and vdW heterostructures 22, but never before at an FM-AF interface.
In our demonstration, an AF-ordered CrSBr bilayer is adjacent to a trilayer with nonzero magnetization, and magnetic-field control of the trilayer is used to control the Néel vector in the bilayer.
@@
Figure 1: Controllable lateral exchange bias. a-i Schematics of the exfoliated CrSBr sample after zero-field cooling. A diamond tip with an embedded NV center, and the angle between the NV center and the sample plane are represented. ii-iv Schematics of the sample after the successive application of 140 mT, 160 mT and 0 mT fields along the crystalline b-axis. b Optical image of the flake obtained by measuring the reflectance of a 640 nm laser. Stacks of 2, 3 and 7 layers are labeled. c Dual-Iso-B (see SI,23) magnetic image of the flake for an external in-plane field of $2$ mT.
@@
Figure 1a-i shows a representative cross-section through our sample that consists of a CrSBr bilayer with adjacent flakes of three and seven layers, respectively (see Fig. 1b for an optical image), and the expected spin configuration after zero-field cooling of the sample 24. CrSBr is an easy axis AF, where the easy axis (the “$b$-axis”) lies in the vdW plane and where spins order ferromagnetically within a plane and antiferromagnetically between planes 19.
As a result, CrSBr samples with an even (odd) number of layers carry zero (nonzero) magnetization, respectively.
@@
To assess the magnetic state of our sample, and image its spin textures, we employ direct magnetic imaging using scanning nitrogen-vacancy (NV) magnetometry. In short, NV magnetometry exploits the electronic spin of the NV centre as a sensitive magnetometer that can be initialised and read out optically and driven by microwave magnetic fields 25. To achieve nanoscale imaging, the NV spin is embedded in a diamond tip26 to be scanned within $\approx 50~{}$nm from the sample. In this work, we exploit both qualitative (Dual-Iso-B) and quantitative (ODMR) imaging modes as described in the Supplementary Information (SI), where we use the former for rapid assessment of spin configurations and the latter for detailed quantitative analysis of the resulting spin textures. Figure 1c shows a qualitative NV magnetometry image of our sample that shows stray field patterns consistent with the spin arrangement illustrated in Fig. 1a-i (see SI,23). All data we present have been obtained at a temperature of $T\approx 4~{}$K in a cryogenic scanning NV magnetometry apparatus with vector magnetic field control described elsewhere 27, 24. When performing NV magnetometry in external magnetic fields, we align these fields with the NV spin quantization axis $\mathbf{e}_{\rm NV}$ to ensure optimal performance 28, and mount our samples such that $\mathbf{e}_{\rm NV}$ lies in the sample’s $b$-$c$-plane, tilted $\approx 54~{}^{\circ}$ from the $c$-axis (Fig. 1a-i). Due to the weak response of CrSBr to magnetic fields along $c$29, we state the magnitude of the $b$-axis component, $B_{b}$, whenever a field is applied.
@@
A possible approach to exploit the LEB for Néel vector control is shown in the sequence of illustrations in Fig. 1a, i-iv. Following a sequence of increasing magnetic field $B_{b}$, where first, the bilayer undergoes a spin-flip transition at $B_{b}\approx 140~{}$mT 30, 31, 32, 29, and subsequently, the trilayer switches its magnetisation at $B_{b}\approx 160~{}$mT 5. Importantly, the $7$-layer flake (that we will refer to as “pinning layer” in the following) remains unaffected by this process due to the monotonic increase of the spin-flip field, $B_{\rm SF,n}$, with the number of layers, $n$33.
The flip of the trilayer (that we will refer to as the “control layer” in the following) exposes the bilayer to LEB of opposite signs at the boundaries to the control and pinning layer, respectively. Upon reduction of $B_{b}$ below $B_{\rm SF}$, AF phases of opposing Néel vector orientation will thus emerge from the two boundaries, resulting in the expected final spin arrangement shown in Fig. 1a-iv.
@@
Figure 2: Lateral exchange bias induced phase wall steering. a Schematics of the magnetic state in the flake for an external field of 144144144~{}144mT (Field history in SI 23). b Dual-Iso-B magnetic image of the flake in the region highlighted in the schematics. c Illustration of the phase wall distortion at the pinning layer interface. The boundary part painted in red has a higher interfacial energy than the green part. d ODMR magnetic image of the interface region. e Micromagnetic simulation of the flake following color code in a (Simulation details in SI 23) f-j Same as a-e with the control layer’s magnetization flipped.
We begin by experimentally assessing the effectiveness of the LEB. For this, we focus on the regime $B_{b}\approx B_{\rm SF,2}$, where the bilayer is in a state of phase-coexistence between FM and AF phases 24, 34. Fig. 2a shows a schematic of the magnetic state of the flake that we infer from the magnetic image obtained at $B_{b}=144~{}$mT (Fig. 2b). Importantly, we find that the LEB has a striking effect on the trajectory of AF-FM phase walls (PW) when they impinge on the adjacent pinning layer.
Specifically, towards the intersection with the bilayer-pinning-layer interface, the PWs show striking and reproducible distortions away from the normal incidence observed and expected 35 when the PW impinges on the flake boundary. These distortions are well explained by the energetics of the LEB at the bilayer-pinning-layer interface (see Fig. 2c):
Where the bilayer has FM spin alignment, the interface incurs an energy penalty corresponding to a head-to-head domain wall [^1] Note that this description holds true only if the bilayer-pinning-layer interface runs along the crystal $a$-axis, as is the case here. in a single CrSBr monolayer. Conversely, if the bilayer is in the AF phase, the interface energy depends on its Néel-vector orientation. For the arrangement represented in Fig. 2c, the interface energy corresponds to a head-to-head DW in two CrSBr monolayers, while for the opposite Néel vector orientation, the interface energy is zero.
To minimize energy, the length of the low-energy bilayer-pinning-layer interface will thus expand at the expense of the high-energy interface. This expansion distorts and extends the PW, where both represent additional energy penalties. The process stops when the energy reaches a local minimum, leading to the distorted PW trajectory shown in Fig. 2d. This behaviour can be seen as a magnetic analogy of wetting in hydrostatics, where the PW forms a “contact angle” with the pinning layer (see SI 23).
@@
We confirmed this intuitive picture through numerical micromagnetic simulations described in detail in the SI 23. Figure 2e shows a representative simulation result for the steady state spin configuration in a model CrSBr system that mimics the geometry of our sample. Importantly, the simulation yields a PW trajectory and FM/AF coexistence that are in good qualitative agreement with our data.
@@
Next, we performed a control experiment in which we investigate the effect of LEB on the PW trajectory for the opposite Néel-vector orientation compared to the previous case (Fig. 2f-j). We first applied a positive field $B_{b}=340~{}$mT to invert the control layer magnetization. When reducing $B_{b}$ towards $B_{\rm SF,2}$, an AF pocket preferentially nucleates at the interface to the control layer 24, and, owing to LEB, will have its Néel vector flipped compared to the case discussed before. This assertion is confirmed by the strikingly different PW behavior that we observe in this case at $B_{b}\approx B_{\rm SF,2}$ (Fig. 2g). The PW is now deflected in the opposite deflection compared to before as a result of the modified energetics at the bilayer-pinning-layer interface. This picture is again verified by the micromagnetic model as shown in Fig. 2j.
@@
Figure 3: LEB enabled anti-ferromagnetic domain wall control. a Schematics of the flake magnetic state at low field. b Dual-Iso-B magnetic images of the flake as a function of a decreasing external field. The AF DW is highlighted on the last panel. c, d Same as a, b with the control layer flipped.
@@
We now apply our knowledge of the LEB for the deterministic writing of a DW in our AF-ordered CrSBr bilayer. Starting from the configuration shown in Fig. 2b, we reduce the magnetic field to below $B_{\rm SF,2}$, where all of the bilayer is AF ordered. Figure 3b shows a corresponding sequence of magnetic images that display the evolution of the bilayer spin texture with decreasing field. Strikingly, at $B_{b}=131~{}$mT, we find that the fully AF ordered bilayer is divided by a well-pronounced line of nonzero stray field. Based on our assessment of the bilayer’s spin structure (Fig. 2f) we associate this line with the presence of an AF DW. A control experiment further confirms this interpretation: If the sequence is repeated from the starting configuration presented in Fig. 2e, f no domain-wall would be expected, and indeed, a perfectly homogeneous AF region results in the bilayer (Fig. 3d).
Figure 4: Decoration of antiferromagnetic order. a Dual-iso-B magnetic image of AFM DW intersecting a thin FM layer and imprinting onto it. b, c ODMR magnetic images of the AFM-FM DW interface at the initialisation field of $B_{b}=138~{}$mT and with the field reduced to near zero $B_{b}=2~{}$mT, respectively. d Schematic of the in-plane spin rotations of each layer at the DW interface. e Linecuts of the magnetic field from the bilayer (blue) and trilayer (green) DWs, as shown in a and b, with an approximate 25$\times$ difference in magnitude.
@@
While a determination of the exact DW structure is out of the scope of this work, our simulations indicate that it is of Néel nature 23. The observed, nonzero DW stray field appears to originate from spin-canting induced by the in-plane ($b$-axis) component of the applied field, as its magnitude depends on the DW angle and increases with the field.
When repeating the experiment, we observed the DW at different, random locations on the sample (see SI 23), which indicates that domain-wall motion is not strongly affected by pinning. We also successfully applied this sequence and produced a DW on an additional bilayer sample (See SI 23).
@@
Finally, our experiment also revealed an instance of AF domain decoration. Figure 4a shows the outcome of a new instance of the domain-wall writing sequence, where the resulting DW crosses a $\approx 200~{}$nm wide CrSBr trilayer that extends from the control layer into the bilayer (see SI 23). A quantitative magnetometry scan (Fig. 4b and c) reveals that, by virtue of the AF interlayer exchange interaction, the domain structure of the bilayer is imprinted onto that trilayer that is now split into a two-domain state (see illustration in Fig. 4d). The resulting twenty fold increase of the DW stray field on the narrow trilayer compared to the bulk bilayer (Fig. 4e), suggests that such thin stripes act as efficient decorations for underlying AF domains and may thereby enable detection of such domains near zero magnetic field, where the DW stray-field is otherwise undetectable.
In conclusion, we have established the LEB as a novel tool for Néel-vector control in vdW AFs. Importantly, our concept readily extents to other vdW magnets and should, in particular, apply to any $a$-type vdW AF, including the prominent examples CrI32, CrCl337or CrPS438, where our “wetting” angle methodology could also be applied to assess interfacial exchange energies.
@@
Our results build on the single crystalline nature of vdW magnets, which offers atomically sharp lateral interfaces — a key factor that sets them apart from their thin-film counterparts and that results in a remarkable enhancement in the reach of EB: While in thin-film geometries 11, 12, each interfacial spin from the pinning layer controls tens of spins in typically nanometres-thin target layers, in our LEB, each interfacial spin controls a row of spins extending microns into the target layer — an extension of the effective range of EB by several orders of magnitude.
@@
Our work opens up exciting future avenues, not only towards the fundamental understanding of DWs 39 and domain formation in atomically thin AFs, but also in the recently explored interplay between magnetism and optical and magnetic excitations of CrSBr 5. Indeed, excitons and magnons appear strongly coupled in this system 40, which, together with magnon guiding on spin textures 41, might enable engineered magnon-exciton dynamics, controlled by LEB-written AF DWs.

### v2

@@
Atomically thin van der Waals (vdW) magnets have emerged as a fascinating platform for the exploration of novel physical phenomena arising from their reduced dimensionality and exceptional material properties. Their single-crystalline nature and ultimate miniaturization position them as leading candidates for next-generation spintronic applications. Antiferromagnetic (AF) vdW magnets are of particular interest, as they combine the advantages of vdW magnets with the functionality of AF spintronics, offering unique opportunities for ultrafast and robust spintronic devices. However, the lack of approaches to locally and deterministically manipulate their order parameter – the Néel-vector – remains a key limitation. Here, we introduce a fundamentally new paradigm in nanomagnetism, which we term lateral exchange bias (LEB), to achieve Néel vector control in bilayers of the vdW AF CrSBr. We exploit the single-crystalline registry formed by terraced CrSBr samples, where the bilayer Néel vector is controlled by LEB from neighboring, odd-layered flakes, whose nonzero magnetization we manipulate using magnetic fields. Using this control, we achieve nonvolatile manipulation of magnetic domains and domain walls in AF CrSBr bilayers, establishing a powerful toolkit for controlling atomically thin AFs at the nanoscale. Our results challenge conventional views on exchange bias and provide a previously unexplored mechanism for achieving atomic-scale control of AFic order. Our findings pave the way for the development of advanced spintronic architectures and quantum technologies based on vdW magnets.
@@
The order parameter of few-layer vdW magnets can be controlled by a range of external stimuli, including magnetic 1, 2 and electric fields 3, 4, doping 4, 5, pressure 6, and strain 7. However, current approaches do not provide the sort of local control over order parameters that is required to deterministically create spin textures such as magnetic domain walls, and, in most cases, do not allow bistable switching between nonvolatile states which is a prerequisite for vdW spintronics. While important first steps have been taken in this direction 8, realizing such local control in fully compensated vdW AFs remains challenging, and so far no controlled writing of spin textures has been demonstrated.
@@
In classical magnetism 9, 10 and thin-film spintronics devices 11, 12, the exchange bias (EB) phenomenon has been known for decades as an efficient approach to controlling magnetic order parameter, and was instrumental for the development of memory and sensor technologies 13. The EB effect is based on interfacial exchange interactions between two adjacent magnetic layers, where the order parameter of a pinning layer, typically an AF, controls the order parameter in the adjacent target layer, typically a ferromagnet (FM). Despite intense research on EB in vdW magnets 14, 15, 16, 17, its use for vdW spintronics and order-parameter control in vdW magnets remains challenging due to the weakness of the interlayer EB and difficulties in reproducibly controlling the vdW gap in vertical vdW heterostructures 18.
@@
Here, we introduce a novel concept of EB in vdW AFs that is intrinsically free of these limitations and demonstrate its use for local control of the order parameter and spin-texture initialization in bilayers of the vdW AF CrSBr 19. Our concept, which we term lateral exchange bias (LEB), exploits exchange interactions across the one-dimensional interface between two adjacent regions with different numbers of layers of the same single-crystal vdW magnet for order-parameter control. Lateral exchange has previously been observed with FM thin films 20, 21 and vdW heterostructures 22, 23, but never before to manipulate the AF order. In our demonstration, an AF-ordered CrSBr bilayer is adjacent to a trilayer with nonzero magnetization, and magnetic-field control of the trilayer is used to control the Néel vector in the bilayer.
@@
Figure 1: Controllable lateral exchange bias. a-i Schematics of the exfoliated CrSBr sample after zero-field cooling. A diamond tip with an embedded NV center, and the angle between the NV center and the sample plane are represented. $i$-iv Schematics of the sample after the successive application of 140 mT, 200 mT and 0 mT fields along the crystalline b-axis. b Optical image of the flake obtained by measuring the reflectance of a 640 nm laser. Stacks of 2, 3 and 7 layers are labeled. c Dual-Iso-B (see SI section III 24) magnetic image of the flake for an external in-plane field of 22 mT.
@@
Figure 1a-i shows a representative cross-section through our sample that consists of a CrSBr bilayer with adjacent flakes of three and seven layers, respectively (see Fig. 1b for an optical image), and the expected spin configuration after zero-field cooling of the sample 25. CrSBr is an easy axis AF, where the easy axis (the “$b$-axis”) lies in the vdW plane and where spins order ferromagnetically within a plane and antiferromagnetically between planes 19. As a result, CrSBr samples with an even (odd) number of layers carry zero (nonzero) magnetization, respectively.
@@
To assess the magnetic state of our sample, and image its spin textures, we employ direct magnetic imaging using scanning nitrogen-vacancy (NV) magnetometry. In short, NV magnetometry exploits the electronic spin of the NV centre as a sensitive magnetometer that can be initialised and read out optically and driven by microwave magnetic fields 26. To achieve nanoscale imaging, the NV spin is embedded in a diamond tip27 to be scanned within $\approx 50$~nm from the sample. In this work, we exploit both qualitative (Dual-Iso-B) and quantitative (ODMR) imaging modes as described in the Supplementary Information (SI) section III 24, where we use the former for rapid assessment of spin configurations and the latter for detailed quantitative analysis of the resulting spin textures. Figure 1c shows a qualitative NV magnetometry image of our sample that shows stray field patterns consistent with the spin arrangement illustrated in Fig. 1a-i 25. All data we present have been obtained at a temperature of $T\approx 4$~K in a cryogenic scanning NV magnetometry apparatus with vector magnetic field control described elsewhere 28, 25. When performing NV magnetometry in external magnetic fields, we align these fields with the NV spin quantization axis $\mathbf{e}_{\rm NV}$ to ensure optimal performance 29, and mount our samples such that $\mathbf{e}_{\rm NV}$ lies in the sample’s $b$-$c$-plane, tilted $\approx 54$$^{\circ}$ from the $c$-axis (Fig. 1a-i). Due to the weak response of CrSBr to magnetic fields along cc30, we state the magnitude of the $b$-axis component, $B_{b}$, whenever a field is applied.
@@
A possible approach to exploit the LEB for Néel vector control is shown in the sequence of illustrations in Fig. 1a, i-iv. Following a sequence of increasing magnetic field $B_{b}$, where first, the bilayer undergoes a spin-flip transition at $B_{b}\approx 140$~mT 31, 32, 33, 30, and subsequently, the trilayer switches its magnetisation at $B_{b}\approx 200$~mT 5. Importantly, the $7$-layer flake (that we will refer to as “pinning layer” in the following) remains unaffected by this process as the spin-flip field values in few layers CrSBr depend on the exact number of layers, and tend to increase with the flake thickness 34 (see SI section VI 24). The flip of the trilayer (that we will refer to as the “control layer” in the following) exposes the bilayer to LEB of opposite signs at the boundaries to the control and pinning layer, respectively. Upon reduction of $B_{b}$ below $B_{\rm SF}$, AF phases of opposing Néel vector orientation will thus emerge from the two boundaries, resulting in the expected final spin arrangement shown in Fig. 1a-iv.
@@
Figure 2: Lateral exchange bias induced phase wall steering. a External magnetic field sequence applied along the easy axis of the flake, with expected magnetization for each stack. b-f correspond to step 4 and g-k to step 2. See SI section VI and VII for more details 24. b Schematics of the magnetic state in the flake for an external field of 144144~mT. c Dual-Iso-B magnetic image of the flake in the region highlighted in the schematics. See SI section IV 24 for a full scale imaging of the bilayer. Dotted lines indicate the boundaries between regions of different thicknesses. d Illustration of the phase wall distortion at the pinning layer interface. The boundary part painted in red has a higher interfacial energy than the green part. Deviation from normal incidence is noted by a black arrow. e ODMR magnetic image of the interface region. Dotted line and black arrow indicate deviation from normal incidence. f Micromagnetic simulation of the flake following color code in a (Simulation details in SI section V 24) g-k Same as b-f with the control layer’s magnetization flipped.
We begin by experimentally assessing the effectiveness of the LEB. For this, we focus on the regime $B_{b}\approx B_{\rm SF,2}$, where the bilayer is in a state of phase-coexistence between FM and AF phases 25, 35, corresponding to the steps 2 and 4 of the external field sequence presented in Fig. 2a. Fig. 2b shows a schematic of the magnetic state of the flake that we infer from the magnetic image obtained at $B_{b}=144$~mT (Fig. 2c). Importantly, we find that the LEB has a striking effect on the trajectory of AF-FM phase walls within the bilayer when they impinge on the adjacent pinning layer. Specifically, towards the intersection with the bilayer-pinning-layer interface, the phase walls show striking and reproducible distortions away from the normal incidence observed and expected 36 when the phase wall impinges on the flake boundary. These distortions are well explained by the energetics of the LEB at the bilayer-pinning-layer interface (see Fig. 2d): Where the bilayer has FM spin alignment, the interface incurs an energy penalty corresponding to a head-to-head domain wall [^1] Note that this description holds true only if the bilayer-pinning-layer interface runs along the crystal $a$-axis, as is the case here. in a single CrSBr monolayer. Conversely, if the bilayer is in the AF phase, the interface energy depends on its Néel-vector orientation. For the arrangement represented in Fig. 2d, the interface energy corresponds to a head-to-head domain wall in two CrSBr monolayers, while for the opposite Néel vector orientation, the interface energy is zero. To minimize energy, the length of the low-energy bilayer-pinning-layer interface will thus expand at the expense of the high-energy interface. This expansion distorts and extends the phase wall, where both represent additional energy penalties. The process stops when the energy reaches a local minimum, leading to the distorted phase wall trajectory shown in Fig. 2e. This behaviour can be seen as a magnetic analogy of wetting in hydrostatics, where the phase wall forms a “contact angle” with the pinning layer (see SI section VIII 24).
@@
We confirmed this intuitive picture through numerical micromagnetic simulations described in detail in the SI section V 24. Figure 2e shows a representative simulation result for the steady state spin configuration in a model CrSBr system that mimics the geometry of our sample. Importantly, the simulation yields a phase wall trajectory and FM/AF coexistence that are in good qualitative agreement with our data.
@@
Next, we performed a control experiment in which we investigate the effect of LEB on the phase wall trajectory for the opposite Néel-vector orientation compared to the previous case (Fig. 2g-k). We first applied a positive field $B_{b}=340$~mT to invert the control layer magnetization. When reducing $B_{b}$ towards $B_{\rm SF,2}$, an AF pocket preferentially nucleates at the interface to the control layer 25, and, owing to LEB, will have its Néel vector flipped compared to the case discussed before. This assertion is confirmed by the strikingly different phase wall behavior that we observe in this case at $B_{b}\approx B_{\rm SF,2}$ (Fig. 2g). The phase wall is now deflected in the opposite deflection compared to before as a result of the modified energetics at the bilayer-pinning-layer interface. This picture is again verified by the micromagnetic model as shown in Fig. 2j.
@@
Figure 3: LEB enabled anti-ferromagnetic domain wall control. a Schematics of the flake magnetic state at B $\approx$ 140 mT when the control layer is flipped to the left. b Dual-Iso-B magnetic images of the flake as a function of a decreasing external field. Dotted lines indicate the boundaries between regions of different thickness. c Schematics of the flake magnetic state at B $\approx$ 130 mT. d-e Same as a-c with the control layer flipped to the right.
@@
We now apply our knowledge of the LEB for the deterministic writing of a domain wall in our AF-ordered CrSBr bilayer. Starting from the experimental state shown in Fig. 2b, we reduce the magnetic field to below $B_{\rm SF,2}$, where all of the bilayer is AF ordered. Figure 3b shows a corresponding sequence of magnetic images that display the evolution of the bilayer spin texture with decreasing field. Strikingly, at $B_{b}=131$~mT, we find that the bilayer is divided by a well-pronounced line of nonzero stray field. Based on our assessment of the bilayer’s spin structure we associate this line with the presence of an AF domain wall (Fig. 3c). A control experiment further confirms this interpretation: If the sequence is repeated from the starting configuration presented in Fig. 3d no domain-wall would be expected, and indeed, a perfectly homogeneous AF region results in the bilayer (Fig. 3e).
[…]
@@
While a determination of the exact domain wall structure is out of the scope of this work, our simulations indicate that it is of Néel nature. The observed, nonzero domain wall stray field appears to originate from spin-canting induced by the in-plane ($b$-axis) component of the applied field, as its magnitude depends on the domain wall angle (excluding an out of plane canting) and increases with the applied field. When repeating the experiment, we observed the domain wall at different, random locations on the sample (see SI section IX 24), which indicates that domain-wall motion is not strongly affected by pinning. We also successfully applied this sequence and produced a domain wall on an additional bilayer sample (See SI section X 24).
@@
Finally, our experiment also revealed an instance of AF domain decoration. Figure 4a shows the outcome of a new instance of the domain-wall writing sequence, where the resulting domain wall crosses a $\approx 200$~nm wide CrSBr trilayer that extends from the control layer into the bilayer (see SI section I 24). A quantitative magnetometry scan (Fig. 4b and c) reveals that, by virtue of the AF interlayer exchange interaction, the domain structure of the bilayer is imprinted onto that trilayer that is now split into a two-domain state (see illustration in Fig. 4d). The resulting twenty fold increase of the domain wall stray field on the narrow trilayer compared to the bulk bilayer (Fig. 4e), suggests that such thin stripes act as efficient decorations for underlying AF domains and may thereby enable detection of such domains near zero magnetic field, where the domain wall stray-field is otherwise undetectable.
@@
In conclusion, we have established the LEB as a novel tool for Néel-vector control in vdW AFs. Importantly, our concept readily extents to other vdW magnets and should, in particular, apply to any $a$-type vdW AF, including the prominent examples CrI32, CrCl338or CrPS439, where our “wetting” angle methodology could also be applied to assess interfacial exchange energies. Our results build on the single crystalline nature of vdW magnets, which offers atomically sharp lateral interfaces — a key factor that sets them apart from their thin-film counterparts and that results in a remarkable enhancement in the reach of EB: While in thin-film geometries 11, 12, each interfacial spin from the pinning layer controls tens of spins in typically nanometres-thin target layers, in our LEB, each interfacial spin controls a row of spins extending microns into the target layer — an extension of the effective range of EB by several orders of magnitude.
@@
Our work opens up exciting future avenues, not only towards the fundamental understanding of domain walls 40 and domain formation in atomically thin AFs, but also in the recently explored interplay between magnetism and optical and magnetic excitations of CrSBr 5. Indeed, excitons and magnons appear strongly coupled in this system 41, which, together with magnon guiding on spin textures 42, might enable engineered magnon-exciton dynamics, controlled by LEB-written AF domain walls.